OpenStudy (anonymous):

Anyone help me with this please?: Making the simplistic assumption that the dissolved NaCl(s) does not affect the volume of the solvent water, determine the constants m and b in the equation Molarity = m*density + b that relates the NaCl molarity to the NaCl(aq) density. Take the density of water to be 1.00 g/ml and the molar mass of NaCl to be 58.5 g/mol.

OpenStudy (goformit100):

@aaronq Please can you Help here ?

OpenStudy (aaronq):

Hm, i'm sute you have: \(M_{NaCl_(aq)}=x*\rho_{NaCl(aq)}+y\) you can substitute: \(\dfrac{n_{solute}}{L_{solution}}=x*\dfrac{m_{solute}}{V_{solution}}+y\) substitute further: \(\dfrac{\dfrac{m_{NaCl}}{M_{NaCl}}}{L_{solution}}=x*\dfrac{m_{NaCl}}{V_{solution}}+y\) You could assume that \(L_{solution}=V_{solution}=1\;L\) then i would graph it

OpenStudy (aaronq):

and so you have:***

OpenStudy (aaronq):

or you can solve for x and y algebraically, if you notice, it's the equation of a line y=mx+b m=\(\dfrac{y_2-y_1}{x_2-x_1} =\dfrac{\dfrac{m_2}{M_{NaCl}}-\dfrac{m_1}{M_{NaCl}}}{m_2-m_1}\) then solve for b like you would for any old line equation

OpenStudy (anonymous):

Sorry, was away for a bit...Thank you goformit100 and thank you aaronq!

OpenStudy (aaronq):

no problem dude!